OpenStudy (anonymous):

solve each system of equations algebraically. 2x-y=8 x^2 + y^2 = 9 I tried solving for y in the first equation, but when I put it in the other equation, I messed something up, but I can't figure out what.

OpenStudy (anonymous):

y = 2x - 8 = x - 4 . Replace this value in the second equation: x^2 + (x - 4)^2 - 9 = 0 2x^2 - 8x + 7 = 0. Use the quadratic equation to solve this equation. D = 64 - 56 = 8 = 4*2 -> VD = 2V2 x1 = 8/4 + 2V2/4 = 2 + V2/2, and x2 = 8/4 - 2V2/4 = 2 - V2/2